Express as a single logarithm:
3log6N - ((3log6 N) / 2)

Answer by venugopalramana(3286)

3log6N - ((3log6 N) / 2)=log6(N^3) - ((log6 (N^3) / 2)
=(log6(N^3))/2=log6((N^3)^(1/2))=log6(N^3/2)